Directrix (directrix):

Four students took a standardized mathematics test and earned scores of a, b, c, and d, with a ≤ b ≤ c ≤ d. The mean of their scores was 645, the median was 640, and the range was 70. What is the largest possible value of c?

OpenStudy (anonymous):

\[{{b+c} \over 2 }= 640\]\[d-a=70\]\[{{a+b+c+d}\over4}=645\]since we know b + c we can reduce the third equation to \[{{a+1280+d}\over4}=645\]now you have 2 equations and two unknowns and you can solve for a and d. This gives you the upper and lower limit. Try it from there. I'm not saying this is the absolute correct way to go about it,but this is where I'd start

OpenStudy (mathmath333):

by the chme solutions c is coming \(\Large c_{max }=665\)